Magical Operations in Lords 32

Ryan Powers

August 22, 2001

1 Basic Concepts of Magical Operations

Magical Operations are magical feats beyond the scale of the simpler, more personal spells available to leaders. Things that modify entire regions, armies, or even nations themselves are magical operations.

Operations must be researched. For more information on researching magical operations, actions, and spells see the Magical Research document.

Each Operation requires one MOC (Magical Operations Capacity) point and may require one or more MOB (Magical Operations Bonus) points. Some may also improved with the expenditure of MFP (Magical Force Points), though none require them.

Magical Operations may be used in any regions controlled by the nation, or up to their Action Range away. (Just like Espionage actions) If a leader is providing the MOC point, the operation must occur in the same region or city as the leader.

Each Operations is presented in a format like this:

1.1 Operation Name

Code The letter code for this operation.

Magic Level The level that be reached in the region(s) that the operation is being used in.

Magical Tech Level The minimum Magical Tech Level required to learn this operation.

MOB The number of Bonus points required.

Results What happens when this operation is used.

2 Common Magical Operations

All nations have knowledge of the existence of the following operations. They still must be researched to be used. A few nations will start out with access to one or more of these, this will be noted on their stat sheet.

2.1 Bountiful Harvest

Code M-BH

Magic Level 1

Magic Tech Level 1

MOB 1 per region affected

Results This operation affects one or more regions. Each region produces more Agro points than it normally would. The amount of the increase is based on how successful the operation is. MFP may be added to increase the success rate and level, or can be used to sustain this operation over multiple turns at a rate of 2 MFP / region / turn.

2.2 Prosperity

Code M-P

Magic Level 1

Magic Tech Level 1

MOB 1 per region affected

Results Same as for Bountiful Harvest except that GP production is altered. Cna also be applied to cities, with each 3GPv of a city counting as one region for the purpose of costs.

2.3 Magical Fair

Code M-MF Magic Level 1 Magic Tech Level 1 MOB 1 for every region beyond the first. Results This operation provides additional MFP form a region

or city. The first region is free of MOB point cost. Each extra costs 1 MOB. For the purposes of costs a city counts as one region per 3GPv rounded up. This operations represents public displays of magic as well as magical contests. The purpose being to recruit more people and thereby increase MFP production for the turn.

2.4 Iron Will

Code M-IW Magic Level 1 Magic Tech Level 1 MOB 1 for every army after the first. Results This operation boosts the morale of a single army with-

out a cost in MOB points. Additional armies may be affected at the same time at a cost of 1 MOB per additional army. This may be maintained additional turns at the rate of 2 MFP / army / turn.

2.5 Fear

Code M-F Magic Level 1 Magic Tech Level 1 MOB 1 per army affected. Results This operation reduces the morale of an army or armies.

It affects them for the entire turn. The chance of success is greatly reduced unless the location of the army is known.

2.6 Dispel Operation

Code M-DO Magic Level 1 Magic Tech Level 1 MOB 1 / target Results This operation attempts to dispel all magical opera-

tions affecting a given target (region, city, army). Each operation being dispelled is tested individually, so partial success is possible. For the purpose of this operation a city is one target regardless of its size.